For a hypothetical chemical reaction that has the stoichiometry 2X + Y -----Z the following initial data were obtained

X (Mol/L Y(Mol/L) rate of formation of Z (Mol/L s

.4 .2 1.4 x 10^-3

.4 .4 2.8 x 10^-3

.2 .1 7.0 x 10^-4

.6 .6 4.2 x 10^-3

Give rate Law for this reaction

Callculate rate constant and specify units

How long must reaction proceed to produce concentration of Z =to .2 molar if initial concentrations of x=.8M and Y = .6 M and Z = 0
Answer #1

Let the rate law of the reaction is given by

Rate = k[X]^a[Y]^b

using the data given in the table we get

1.4 * 10^{-3} = k[0.4]^a[0.2]^b --- (i)

2.8 * 10^{-3} = k[0.4]^a[0.4]^b --- (ii)

dividing the first equation by second equation we get

2^b = 2=> b=1

2^a * 2 = 2 => a = 0

Hence the rate law of the equation will be given

Rate = k[Y]

rate constant = 7 * 10^(-3) s^(-1)

ln(Ao/At) = kt

ln(0.6/0.4) = kt

t = ln(1.5)1000/7 = 57.923 seconds